### v2

@@ Abstract
Speech-driven 3D talking heads generation has emerged as a significant area of interest among researchers, presenting numerous challenges. Existing methods are constrained by animating faces with fixed topologies, wherein point-wise correspondence is established, and the number and order of points remains consistent across all identities the model can animate. In this work, we present ScanTalk, a novel framework capable of animating 3D faces in arbitrary topologies including scanned data. Our approach relies on the DiffusionNet architecture to overcome the fixed topology constraint, offering promising avenues for more flexible and realistic 3D animations. By leveraging the power of DiffusionNet, ScanTalk not only adapts to diverse facial structures but also maintains fidelity when dealing with scanned data, thereby enhancing the authenticity and versatility of generated 3D talking heads. Through comprehensive comparisons with state-of-the-art methods, we validate the efficacy of our approach, demonstrating its capacity to generate realistic talking heads comparable to existing techniques. While our primary objective is to develop a generic method free from topological constraints, all state-of-the-art methodologies are bound by such limitations. Code for reproducing our results, and the pre-trained model will be made available.
@@ 1 Introduction
Human face animation is a complex task, widely explored in Computer Vision and Computer Graphics because of the broad range of applications drawn from it, spanning from virtual reality to video game graphics, and more. As 3D face models continue to improve, it becomes more and more relevant to incorporate multi-modal aspects such as speech with audio data. Speech-driven facial animation encounters challenges inherent to the fact that finding a cross-modality mapping from audio to geometric data is an ill-posed problem, dealing with the complex geometry of human faces and the limited availability of paired audio-visual 3D data. Another, less discussed aspect that touches a broader range of 3D graphics with meshes is the robustness to changes in the topology of the mesh, which refers to the arrangement of the vertices and how they are connected. Maintaining fidelity and coherence across different topologies is crucial for ensuring that facial animations remain realistic and expressive, regardless of variations in the underlying mesh structure. This challenge becomes particularly pronounced in speech-driven facial animation, where the dynamics of speech-related lip movements, and related face changes necessitate a high degree of adaptability within the mesh topology.
@@ 2 Related Work
In the literature, there are several methodologies to acquire 3D face meshes, either by extracting the geometry from images or videos [50, 27] or by using a complex set of instruments in a controlled environment [37, 18, 51, 35, 16, 47]. The former offers advantages in terms of easier and more cost-effective data acquisition. However, these methods may sometimes fall short in capturing the complete 3D information from 2D data. With 3D scans, other challenges arise: they come unregistered and may present alterations such as holes and noise. Then, animating these objects directly becomes even more challenging. One can do it by registering the meshes onto a pre-defined topology [24, 29, 47, 16, 23] before animating the registrations with 3D morphable models [14] for example.
@@
Several deep learning models [26, 3, 9, 4] proposed ways to learn a latent representation of face scans using robust encoders such as PointNet [5, 33] but the resulting mesh is registered, which tends to smooth out some details from the scan geometry. However, this extra registration step may be handled efficiently with recent industrial applications such as MetaHuman from Epic Games***https://dev.epicgames.com/documentation/en-us/metahuman/metahuman-documentation. More recently, and closer to our goal, DiffusionNet [39] was combined with neural Jacobian fields [1] in Neural Face Rigging (NFR) [34] to learn a per-triangle deformation field on faces, allowing to transfer an animation from a sequence of unregistered face meshes to another. We stand out from this work by proposing a model that uses audio data to animate a given unregistered face mesh. While DiffusionNet has demonstrated its capability to generalize across varying triangulations in a static setting, our model is the first to exhibit similar properties in a multi-modal 4D setting.
@@
In recent years, numerous models and methodologies have emerged to address the challenge of synchronizing facial animations with speech audio. While significant progress has been made on 2D talking heads [21, 6, 46, 2, 12, 53, 45], only a handful of approaches can be seamlessly extended to 3D data. Procedural techniques have been proposed to animate 3D faces [13, 28, 8, 44, 49], primarily relying on visemes (groups of phonemes) to drive the movements of facial muscles. However, many of these models necessitate re-targeting and struggle to generalize effectively without extensive processing steps. More recent works leverage the increasing availability of data to develop statistical methods, including deep learning strategies [10, 36, 15, 19, 43, 42, 30, 48, 40]. Currently, state-of-the-art deep learning approaches are limited to a fixed mesh topology, requiring it to be identical to the one observed during the training phase. Among these models, VOCA [10] pioneers the development of deep models trained on a large-scale dataset of registered meshes. It was outperformed when MeshTalk [36] introduced a larger registered dataset named Multiface [47] along with a more expressive model. Moving forward, FaceFormer [15] utilized the transformer architecture, optimized for temporal data, and leveraged the power of a large-scale pre-trained audio encoder called Wav2vec2 [38]. This strategy saw further development with CodeTalker [48] and SelfTalk [31]. More recently, FaceXHubert [19] and FaceDiffuser [40] utilized an improved audio encoder, Hubert [20], demonstrating superiority to predict cross-modality mappings from audio to face motions. Speech is also highly correlated with expressions and head pose, as well as eye blink. This motivated recent work to integrate emotions [32, 11] and head poses [52, 41] but the very limited amount of registered data makes the task even more challenging. While continuously enhancing performances, these models are bound to a fixed mesh topology, hindering real-world applications and limiting the quantity of usable data for a single model, both for training and inference.
@@ 3 Proposed approach: ScanTalk
The implementation details of our architecture are provided in the supplementary material for further reference Sec. 6.3.
@@ 3.1 ScanTalk Encoder
Critical to DiffusionNet’s functionality are precomputed features of the face surface, namely the Cotangent Laplacian, Eigenbasis, Mass Matrix, and Spatial Gradient Matrix. These features capture geometric properties of the face, and facilitate information propagation across the surface domain. The Cotangent Laplacian, a discrete version of the Laplace-Beltrami operator, quantifies surface curvature and smoothness during the diffusion process. The Eigenbasis comprises eigenvectors derived from the Laplacian matrix, representing fundamental modes of variation on the surface. The Mass Matrix characterizes the distribution of mass or area across vertices, while the Spatial Gradient Matrix captures spatial derivatives of scalar functions defined on the surface. Together, these operators enhance the architecture’s robustness, flexibility, and expressiveness, making DiffusionNet suitable for diverse surface learning tasks. Notably, this architecture accommodates 3D faces of any topology, allowing for variations in the number and order of defining points.
Let represent the precomputed features obtained by applying the above surface closed-form operators to the 3D neutral face . The DiffusionNet Encoder , with latent size , is designed to process a neutral 3D face mesh , and requires the precomputed per-vertex features to extract a per-vertex descriptors , that is:
@@ 3.2 ScanTalk Decoder
We combine per-vertex descriptors extracted from the neutral face, with the latent vector extracted from the Bidirectional-LSTM:
@@ 4.1 Datasets
For training and quantitative evaluations, we rely on three classical datasets in 3D speech-driven motion synthesis: VOCA [10], BIWI [16] and Multiface [47].
@@
BIWI comprises 14 subjects articulating 40 sentences each, sampled at 25fps. Each sentence lasts around 5 seconds, and the meshes are registered to a fixed topology. Due to GPU limitations, we used a downsampled version of this dataset, called BIWI, which has a fixed topology, with 3,895 vertices and 7,539 faces.
@@
For training purposes, the meshes in both Multiface and BIWI have been scaled and aligned with the meshes in the VOCAset dataset. The data splits can be found in the supplementary material Sec. 6.4. Since ScanTalk is the first topology-independent 3D talking head generator, for the sake of a comprehensive comparison with the state-of-the-art, we trained 4 different models: one is trained using all the three datasets together (multi-dataset), while the other three are trained on each dataset separately (single-dataset).
@@ 4.3 Quantitative Results
Although ScanTalk is the first deep learning model to animate unregistered 3D face meshes, comparisons with the state-of-the-art is possible when inferring our model on registered data. In Sec. 4.3, we present the performance evaluation of ScanTalk training on both a single-dataset and multi-datasets, juxtaposed against several state-of-the-art methods trained solely on a single dataset. While the augmentation of training data enhances the effectiveness of our model, addressing disparities in data composition becomes important for animating unseen faces. Particularly, the varied geometries and head dynamics inherent in BIWI, VOCA, and Multiface meshes pose significant challenges, potentially limiting performance compared to a model trained exclusively on a single dataset.
| VOCAset | BIWI | Multiface | |||||||
| LVE | MVE | FDD | LVE | MVE | FDD | LVE | MVE | FDD | |
| VOCA | 6.993 | 0.983 | 2.662 | 5.743 | 2.591 | 41.482 | 4.923 | 2.761 | 55.781 |
| FaceFormer | 6.123 | 0.935 | 2.163 | 4.085 | 2.163 | 37.091 | 2.451 | 1.453 | 20.239 |
| SelfTalk | 5.618 | 0.918 | 2.321 | 3.628 | 2.062 | 35.470 | 2.281 | 1.901 | 37.434 |
| CodeTalker | 3.549 | 0.888 | 2.258 | 5.190 | 2.641 | 20.599 | 4.091 | 2.382 | 47.905 |
| FaceDiffuser | 4.350 | 0.901 | 2.437 | 4.022 | 2.128 | 39.604 | 3.555 | 2.388 | 29.157 |
| ScanTalk m-d | 6.375 | 0.987 | 2.101 | 4.044 | 2.057 | 40.051 | 2.435 | 1.678 | 32.202 |
@@
This observation becomes evident when assessing Scantalk’s performance on VOCAset, where a substantial improvement is observed when training on a single-dataset. Conversely, in the case of BIWI and Multiface single-dataset training, a decline in results is noted compared to the multi-dataset version. This disparity is attributed to the presence of speech-related head and upper face movements within the ground truth talking head sequences of BIWI and Multiface. Consequently, the multi-dataset training model demonstrates superior performance as it encounters a wider array of sequences featuring head and upper face movements. In contrast, VOCAset lacks such movements, with the 3D faces remaining static, while only the mouth moves, thereby enhancing performance for ScanTalk trained solely on VOCAset. This obviates the need to learn other movements of the head or upper face.
The distinctions between VOCAset and the other two datasets are further underscored by the FDD metric, indicating that 3D faces in VOCAset remain stationary, with minimal activity in the upper facial area and head. Across all tested models, generating sequences with lower FDD is notably more achievable on VOCAset compared to the other datasets, thus emphasizing the disparities between VOCAset and the others. Nevertheless, the results presented in Sec. 4.3 demonstrate that ScanTalk produces outcomes comparable to state-of-the-art methods, whether undergoing single-dataset or multi-dataset training. Moreover, ScanTalk emerges as the first model capable of training in multi-dataset settings, demonstrating the ability to animate a 3D face regardless of its topology.
@@ 4.4 Qualitative Results
In the domain of 3D speech-driven talking heads generation, qualitative evaluations hold as much significance as quantitative assessments. To evaluate the quantitative findings presented in Sec. 4.3, we introduce a straightforward test aimed at discerning disparities in head and upper face movements across the three datasets: for each dataset, we compute the average norm of the differences between the initial frame and subsequent frames within each sequence. This enables the quantification of both head dynamics and upper facial movements across mesh sequences. In the heatmaps of Sec. 4.3, we present the per-frame average difference for each dataset. Our analysis reveals that in the VOCAset, only the lower part of the face exhibits discernible movements, whereas in BIWI and Multiface, the entire head or face moves.
@@
In Fig. 4, 3D faces in different topology, including a scan, animated using ScanTalk are presented. Our method demonstrates good generalization capabilities, successfully animating diverse 3D faces. Notably, enhanced animation quality is observed when a mouth aperture is present; however, our model performs well even in its absence. Nonetheless, without a mouth aperture, the model struggles to generate the corresponding mouth opening, despite accurate lip movement synchronization. More qualitative results can be found in the supplementary material Sec. 6.6.
@@ 4.5 Ablations Studies
In our investigation, as proven by [43, 30, 19], we found that supplementing pre-trained audio encoders with additional temporal consistency mechanisms such as Bidirectional-LSTM, Bidirectional-GRU, or an autoregressive Transformer Decoder (TD) significantly enhances model performance, as illustrated in Tab. 2(b) and in Tab. 3 middle. The BiLSTM model architecture is detailed in Sec. 3, while the BiGRU model mirrors the BiLSTM architecture but substitutes BiLSTM with BiGRU. Details about the Transformer Decoder can be found in the supplementary material Sec. 6.2. From the findings presented in Tab. 2(b) and in Tab. 3 middle, we observe that employing a multilayer Bidirectional-LSTM in the audio stream processing of ScanTalk yields the most favorable performance across both single-dataset training and multi-dataset training scenarios.
@@
| VOCAset | BIWI | Multiface | |||||||
| LVE | MVE | FDD | LVE | MVE | FDD | LVE | MVE | FDD | |
| ST w Wav2Vec2 | 7.127 | 1.241 | 2.349 | 4.439 | 2.276 | 41.374 | 3.198 | 2.329 | 33.671 |
| ST w WavLM | 6.773 | 1.004 | 2.220 | 4.333 | 2.162 | 31.922 | 2.907 | 2.004 | 8.500 |
| ST w TD | 6.512 | 0.995 | 1.858 | 4.827 | 2.168 | 36.789 | 2.136 | 1.725 | 35.044 |
| ST w BiGRU | 6.584 | 0.994 | 2.054 | 4.421 | 2.103 | 41.456 | 2.534 | 1.987 | 34.098 |
| ST | 7.031 | 1.092 | 1.473 | 4.755 | 2.227 | 39.458 | 2.547 | 1.898 | 15.973 |
| ST | 7.451 | 1.084 | 0.859 | 4.748 | 2.199 | 36.939 | 2.888 | 1.891 | 32.501 |
| ST | 6.740 | 0.998 | 1.899 | 4.509 | 2.180 | 33.614 | 2.103 | 1.775 | 31.599 |

### v3

@@ Abstract
Speech-driven 3D talking heads generation has emerged as a significant area of interest among researchers, presenting numerous challenges. Existing methods are constrained by animating faces with fixed topologies, wherein point-wise correspondence is established, and the number and order of points remains consistent across all identities the model can animate. In this work, we present ScanTalk, a novel framework capable of animating 3D faces in arbitrary topologies including scanned data. Our approach relies on the DiffusionNet architecture to overcome the fixed topology constraint, offering promising avenues for more flexible and realistic 3D animations. By leveraging the power of DiffusionNet, ScanTalk not only adapts to diverse facial structures but also maintains fidelity when dealing with scanned data, thereby enhancing the authenticity and versatility of generated 3D talking heads. Through comprehensive comparisons with state-of-the-art methods, we validate the efficacy of our approach, demonstrating its capacity to generate realistic talking heads comparable to existing techniques. While our primary objective is to develop a generic method free from topological constraints, all state-of-the-art methodologies are bound by such limitations. Code for reproducing our results, and the pre-trained model are available at https://github.com/miccunifi/ScanTalk.
@@ 1 Introduction
Human face animation is a complex task, widely explored in Computer Vision and Computer Graphics because of the broad range of applications drawn from it, spanning from virtual reality to video game graphics, and more. As 3D face models continue to improve, it becomes more and more relevant to incorporate multi-modal aspects such as speech with audio data. Speech-driven facial animation encounters challenges inherent to the fact that finding a cross-modality mapping from audio to geometric data is an ill-posed problem, dealing with the complex geometry of human faces and the limited availability of paired audio-visual 3D data. Another, less discussed aspect that touches a broader range of 3D graphics with meshes is the robustness to changes in the topology of the mesh, which refers to the arrangement of the vertices and how they are connected.
Maintaining fidelity and coherence across different topologies is crucial for ensuring that facial animations remain realistic and expressive, regardless of variations in the underlying mesh structure. This challenge becomes particularly pronounced in speech-driven facial animation, where the dynamics of speech-related lip movements, and related face changes necessitate a high degree of adaptability within the mesh topology.
@@ 2 Related Work
In literature, there are several methodologies to acquire 3D face meshes, either by extracting the geometry from images or videos [48, 26] or by using a complex set of instruments in a controlled environment [36, 17, 49, 34, 15, 45]. The former offers advantages in terms of easier and more cost-effective data acquisition. However, these methods may sometimes fall short in capturing the complete 3D information from 2D data. With 3D scans, other challenges arise: they come unregistered and may present alterations such as holes and noise. Then, animating these objects directly becomes even more challenging. One can do it by registering the meshes onto a pre-defined topology [23, 28, 45, 15, 22] before animating the registrations with 3D morphable models [13] for example.
@@
Several deep learning models [25, 3, 9, 4, shape_transformer] proposed ways to learn a latent representation of face scans using robust encoders such as PointNet [5, 32] and Transformers but the resulting mesh is registered, which tends to smooth out some details from the scan geometry. However, this extra registration step may be handled efficiently with recent industrial applications such as MetaHuman from Epic Games. More recently, and closer to our goal, DiffusionNet [38] was combined with neural Jacobian fields [1] in Neural Face Rigging (NFR) [33] to learn a per-triangle deformation field on faces, allowing to transfer an animation from a sequence of unregistered face meshes to another. We stand out from this work by proposing a model that uses audio data to animate a given unregistered face mesh. While DiffusionNet has demonstrated its capability to generalize across varying triangulations in a static setting, our model is the first to exhibit similar properties in a multi-modal 4D setting.
@@
In recent years, numerous models and methodologies have emerged to address the challenge of synchronizing facial animations with speech audio. While significant progress has been made on 2D talking heads [20, 6, 44, 2, 11, 50, 43], only a handful of approaches can be seamlessly extended to 3D data. Procedural techniques have been proposed to animate 3D faces [12, 27, 8, 42, 47], primarily relying on visemes (groups of phonemes) to drive the movements of facial muscles. However, many of these models necessitate re-targeting and struggle to generalize effectively without extensive processing steps. More recent works leverage the increasing availability of data to develop statistical methods, including deep learning strategies [10, 35, 14, 18, 41, 40, 29, 46, 39, 30]. Currently, state-of-the-art deep learning approaches are limited to a fixed mesh topology, requiring it to be identical to the one observed during the training phase. Among these models, VOCA [10] pioneers the development of deep models trained on a large-scale dataset of registered meshes. It was outperformed when MeshTalk [35] introduced a larger registered dataset named Multiface [45] along with a more expressive model. Moving forward, FaceFormer [14] utilized the transformer architecture, optimized for temporal data, and leveraged the power of a large-scale pre-trained audio encoder called Wav2vec2 [37]. This strategy saw further development with CodeTalker [46] and SelfTalk [31]. More recently, FaceXHubert [18] and FaceDiffuser [39] utilized an improved audio encoder, Hubert [19], demonstrating superiority to predict cross-modality mappings from audio to face motions. While continuously enhancing performances, these models are bound to a fixed mesh topology, hindering real-world applications and limiting the quantity of usable data for a single model, both for training and inference.
@@ 3.1 ScanTalk Encoder
Critical to DiffusionNet’s functionality are precomputed features of the face surface, namely the Cotangent Laplacian, Eigenbasis, Mass Matrix, and Spatial Gradient Matrix.
Together, these operators enhance the architecture’s robustness, flexibility, and expressiveness, making DiffusionNet suitable for diverse surface learning tasks. Notably, this architecture accommodates 3D faces of any topology, allowing for variations in the number and order of defining points.
Let represent the precomputed features obtained by applying the above surface closed-form operators to the 3D neutral face .
The DiffusionNet Encoder , with latent size , is designed to process a neutral 3D face mesh , and requires the precomputed per-vertex features to extract a per-vertex descriptors , that is:
@@ 3.2 ScanTalk Decoder
We combine per-vertex descriptors extracted from the neutral face, with the latent vector extracted from the Bidirectional-LSTM
@@ 4.1 Datasets
For training and quantitative evaluations, we rely on three classical datasets in 3D speech-driven motion synthesis: VOCAset [10], BIWI [15] and Multiface [45].
@@
BIWI comprises 14 subjects articulating 40 sentences each, sampled at 25fps.
Each sentence lasts around 5 seconds, and the meshes are registered to a fixed topology. Due to GPU limitations, we used a downsampled version of this dataset, called BIWI6, which has a fixed topology, with 3,895 vertices and 7,539 faces.
@@
For training purposes, the meshes in both Multiface and BIWI6 have been scaled and aligned with the meshes in the VOCAset dataset. The data splits can be found in the supplementary material. Since ScanTalk is the first topology-independent 3D talking head generator, for the sake of a comprehensive comparison with the state-of-the-art, we trained 4 different models: one is trained using all the three datasets together (multi-dataset), while the other three are trained on each dataset separately (single-dataset).
@@ 4.3 Quantitative Results
Although ScanTalk is the first deep learning model to animate unregistered 3D face meshes, comparisons with the state-of-the-art is possible when inferring our model on registered data. In Sec. 4.3, we present the performance evaluation of ScanTalk training on both a single-dataset and multi-datasets, juxtaposed against several state-of-the-art methods trained solely on a single dataset. While the augmentation of training data enhances the effectiveness of our model, addressing disparities in data composition becomes important for animating unseen faces. Particularly, the varied geometries and head dynamics inherent in BIWI6, VOCA, and Multiface meshes pose significant challenges, potentially limiting performance compared to a model trained exclusively on a single dataset.
![[Uncaptioned image]](x3.png)
VOCAset | BIWI6 | Multiface | |||||||||
LVE | MVE | FDD | LVE | MVE | FDD | LVE | MVE | FDD | |||
VOCA | 6.993 | 0.983 | 2.662 | 5.743 | 2.591 | 41.482 | 4.923 | 2.761 | 55.781 | ||
FaceFormer | 6.123 | 0.935 | 2.163 | 4.085 | 2.163 | 37.091 | 2.451 | 1.453 | 20.239 | ||
SelfTalk | 5.618 | 0.918 | 2.321 | 3.628 | 2.062 | 35.470 | 2.281 | 1.901 | 37.434 | ||
CodeTalker | 3.549 | 0.888 | 2.258 | 5.190 | 2.641 | 20.599 | 4.091 | 2.382 | 47.905 | ||
FaceDiffuser | 4.350 | 0.901 | 2.437 | 4.022 | 2.128 | 39.604 | 3.555 | 2.388 | 29.157 | ||
ScanTalk m-d | 6.375 | 0.987 | 2.101 | 4.044 | 2.057 | 40.051 | 2.435 | 1.678 | 32.202 | ||
@@
This observation becomes evident when assessing Scantalk’s performance on VOCAset, where a substantial improvement is observed when training on a single-dataset. Conversely, in the case of BIWI6 and Multiface single-dataset training, a decline in results is noted compared to the multi-dataset version. This disparity is attributed to the presence of speech-related head and upper face movements within the ground truth talking head sequences of BIWI6 and Multiface. Consequently, the multi-dataset training model demonstrates superior performance as it encounters a wider array of sequences featuring head and upper face movements. In contrast, VOCAset lacks such movements, with the 3D faces remaining static, while only the mouth moves, thereby enhancing performance for ScanTalk trained solely on VOCAset. This obviates the need to learn other movements of the head or upper face.
The distinctions between VOCAset and the other two datasets are further underscored by the FDD metric, indicating that 3D faces in VOCAset remain stationary, with minimal activity in the upper facial area and head. Across all tested models, generating sequences with lower FDD is notably more achievable on VOCAset compared to the other datasets, thus emphasizing the disparities between VOCAset and the others.
Nevertheless, the results presented in Sec. 4.3 demonstrate that ScanTalk produces outcomes comparable to state-of-the-art methods, whether undergoing single-dataset or multi-dataset training. Moreover, ScanTalk emerges as the first model capable of training in multi-dataset settings, demonstrating the ability to animate a 3D face regardless of its topology. An interesting characteristic of ScanTalk emerges when analyzing the GPU memory usage relative to the vertex count in the 3D facial model requiring animation. Fig. 3 illustrates a linear increase in GPU usage correlated with the number of vertices.
@@ 4.4 Qualitative Results
In the domain of 3D speech-driven talking heads generation, qualitative evaluations hold as much significance as quantitative assessments. To evaluate the quantitative findings presented in Sec. 4.3, we introduce a straightforward test aimed at discerning disparities in head and upper face movements across the three datasets: for each dataset, we compute the average norm of the differences between the initial frame and subsequent frames within each sequence. This enables the quantification of both head dynamics and upper facial movements across mesh sequences. In the heatmaps of Sec. 4.3, we present the per-frame average difference for each dataset. Our analysis reveals that in the VOCAset, only the lower part of the face exhibits discernible movements, whereas in BIWI6 and Multiface, the entire head or face moves.
@@
In Fig. 5, 3D faces in different topology, including a scan, animated using ScanTalk are presented. Our method demonstrates good generalization capabilities, successfully animating diverse 3D faces. Notably, enhanced animation quality is observed when a mouth aperture is present; however, our model performs well even in its absence. Nonetheless, without a mouth aperture, the model struggles to generate the corresponding mouth opening, despite accurate lip movement synchronization.
@@ 4.5 Ablations Studies
In our investigation, as proven by [41, 29, 18], we found that supplementing pre-trained audio encoders with additional temporal consistency mechanisms such as Bidirectional-LSTM, Bidirectional-GRU, or an autoregressive Transformer Decoder (TD) significantly enhances model performance, as illustrated in Tab. 2(b) and in Tab. 3 middle. The BiLSTM model architecture is detailed in Sec. 3, while the BiGRU model mirrors the BiLSTM architecture but substitutes BiLSTM with BiGRU. Details about the Transformer Decoder can be found in the supplementary material. From the findings presented in Tab. 2(b) and in Tab. 3 middle, we observe that employing a multilayer Bidirectional-LSTM in the audio stream processing of ScanTalk yields the most favorable performance across both single-dataset training and multi-dataset training scenarios.
4.5.3 Loss function.
@@
VOCAset | BIWI6 | Multiface | |||||||||
LVE | MVE | FDD | LVE | MVE | FDD | LVE | MVE | FDD | |||
ST w Wav2Vec2 | 7.127 | 1.241 | 2.349 | 4.439 | 2.276 | 41.374 | 3.198 | 2.329 | 33.671 | ||
ST w WavLM | 6.773 | 1.004 | 2.220 | 4.333 | 2.162 | 31.922 | 2.907 | 2.004 | 8.500 | ||
ST w TD | 6.512 | 0.995 | 1.858 | 4.827 | 2.168 | 36.789 | 2.136 | 1.725 | 35.044 | ||
ST w BiGRU | 6.584 | 0.994 | 2.054 | 4.421 | 2.103 | 41.456 | 2.534 | 1.987 | 34.098 | ||
ST | 7.031 | 1.092 | 1.473 | 4.755 | 2.227 | 39.458 | 2.547 | 1.898 | 15.973 | ||
ST | 7.451 | 1.084 | 0.859 | 4.748 | 2.199 | 36.939 | 2.888 | 1.891 | 32.501 | ||
ST | 6.740 | 0.998 | 1.899 | 4.509 | 2.180 | 33.614 | 2.103 | 1.775 | 31.599 | ||